6-15 Carla Inc. issues $2,100,000 of 7% bonds due in 14 years with interest payable at...

6-15

Carla Inc. issues $2,100,000 of 7% bonds due in 14 years with interest payable at year-end. The current market rate of interest for bonds of similar risk is 9%.

Click here to view factor tables

What amount will Carla receive when it issues the bonds? (For calculation purposes, use 5 decimal places as displayed in the factor table provided and final answer to 0 decimal places, e.g. 458,581.)

Amount received by Carla when bonds were issued $enter a dollar amount rounded to 0 decimal places
Answer #1
Amount received by Carla when bonds were issued 1772989
Workings:
Annual interest 147000 =2100000*7%
Amount PV factor 9% Present value
Annual interest 147000 7.78615 1144564
Principal 2100000 0.29925 628425
Total 1772989
PV factor 9%:
Annual interest 7.78615 =(1-(1.09)^-14)/0.09
Principal 0.29925 =1/1.09^14